# Editable Fairness: Fine-Grained Bias Mitigation in Language Models

### v1

@@ C.4 Debiasing Results on BERT and GPT2
Attribute Race Religion
@@ Appendix D Analysis
D.1 Robustness Analysis of Knowledge Locating
@@
Dataset StereoSet Crows-Pairs WinoBias
Critical Layer 11 11 11
Number of In-domain Samples 537 168 814
Total Number of Samples 771 262 1584
Ratio 69.60% 64.10% 51.40%
@@
D.2 Fine-Tuning vs. Our FAST
@@
Attribute Gender Race
[…]
D.3 Robustness to the Number of Social Biases
We investigate the effectiveness of our proposed method under continual debiasing settings. We perform FAST on different knowledge sets in sequence and evaluate their performance. Results are reported in Table 21. It can be observed that SS obtained in the front stages is steady across the following stages, which indicates that after calibration on other knowledge, existing stored knowledge is still retained. Besides, LMS and ICAT even increase slightly in the process. These results prove the feasibility of continually updating the perception within large language models.
@@ D.3 Robustness to the Number of Social Biases
D.4 Effectiveness of Multi-layer Debiasing
In this section, we discuss the effectiveness of inserting our fairness stamp into multiple layers. We conduct experiments on the located layer and its adjacent one and two layers, with results presented in Table 22 andTable 23. It can be observed that debiasing results vary slightly with the number of layers, illustrating the robustness of one-layer calibration.
@@ D.4 Effectiveness of Multi-layer Debiasing
D.5 Unility Analysis
In Table 23, we report the number of external parameters and time consumed by our proposed method in two stages. The time is counted on a single RTX 3090. The debiasing process can be finished in several seconds, indicating the feasibility of timely LLM calibration.
D.6 Ablation Study on Batch Size
[…]
1 49.87 50.31 47.06 83.33 77.82 77.62
2 51.26 50.94 50.49 73.81 83.04 80.94
4 50.25 52.26 49.57 92.86 84.66 84.24
8 51.17 52.2 49.68 95.24 84.95 82.96
16 53.02 50.94 49.28 92.86 85.39 80.24
32 50.18 55.35 48.5 92.86 85.78 85.47
64 51.34 54.72 50.68 97.62 85.63 83.34
@@ Appendix E Limitation and Future Works
While our research yields important contributions, we acknowledge the presence of certain limitations. Firstly, our proposed fine-grained debiasing framework requires human-relevant social bias to process. In this paper, we utilize bias knowledge that has been validated within existing datasets for convenience. In practice, retaining a comprehensive bias knowledge base is both time-consuming and labor-intensive. We notice that recent works [60, 14] have proposed an automated social bias detection method. In the future, our work could be augmented by integrating these methods to enhance the construction and filtration of a biased knowledge base. Besides, social bias in open language generation or dialogue [74, 55, 19] represents another critical scenario for applying mitigating techniques, which is not addressed in this paper. Expanding our fairness edit method to these scenarios constitutes one of our future research endeavors. Finally, compared to the results on BERT and GPT2, the debiasing performance on larger models (Section 5.2) appears less pronounced. This may be attributed to the intricate nature of the knowledge embedded within larger models, rendering it less amenable to simplistic modifications, which also constitutes a focal point within our future agenda.
Appendix F Broader Impact
With the widespread application of large language models (LLMs), the emphasis on fairness has significantly increased, requiring LLMs to treat individuals from different backgrounds fairly. However, LLMs trained on large datasets inevitably exhibit certain biases during the pre-training phase. In this paper, we propose a promising solution that mitigates unfairness in LLMs while not compromising capability, which is of great significance for deploying fair and reliable language models.

### v2

@@ C.4 Debiasing Results on BERT and GPT2
Dataset StereoSet Crows-Pairs WinoBias
Critical Layer 11 11 11
Number of In-domain Samples 537 168 814
Total Number of Samples 771 262 1584
Ratio 69.60% 64.10% 51.40%
Attribute Gender Race
@@ Appendix D Analysis
D.1 Effectiveness on the Knowledge-editing Task.
We conduct experiments on the knowledge-editing task of Zero-Shot Relation Extraction (zsRE) Levy et al. (2017). We employ GPT-J-6B Wang and Komatsuzaki (2021) as backbone, and use baseline methods including the improved Constrained Fine-Tuning (FT+W) Meng et al. (2022a), MEND Mitchell et al. (2021), ROME Meng et al. (2022a), and MEMIT Meng et al. (2022b). We select layers 3 through 8 for editing, consistent with MEMIT. The training and evaluating datasets are also consistent with MEMIT. As indicated in Table 20, our method outperforms most baseline approaches and achieves performance comparable to MEMIT. These results demonstrate the effectiveness of our method in knowledge-editing tasks. Additional effectiveness validation of fairness stamp (i.e., Section 2.4) is provided in Appendix D.5.
[…]
GPT-J 26.4 (±0.6) 25.8 (±0.5) 27.0 (±0.5)
FT-W 69.6 (±0.6) 64.8 (±0.6) 24.1 (±0.5)
MEND 19.4 (±0.5) 18.6 (±0.5) 22.4 (±0.5)
ROME 21.0 (±0.7) 19.6 (±0.7) 0.9 (±0.1)
MEMIT 96.7 (±0.3) 89.7 (±0.5) 26.6 (±0.5)
FAST 95.1(±0.4) 90.6 (±0.5) 24.6 (±0.5)
D.2 Ablation Study on the Losses.
D.3 Robustness Analysis of Knowledge Localizing
@@
D.4 Ablation Study on Batch Size
[…]
1 49.87 50.31 47.06 83.33 77.82 77.62
2 51.26 50.94 50.49 73.81 83.04 80.94
4 50.25 52.26 49.57 92.86 84.66 84.24
8 51.17 52.2 49.68 95.24 84.95 82.96
16 53.02 50.94 49.28 92.86 85.39 80.24
32 50.18 55.35 48.5 92.86 85.78 85.47
64 51.34 54.72 50.68 97.62 85.63 83.34
@@ D.4 Ablation Study on Batch Size
D.5 Fine-Tuning vs. Our FAST
@@
D.6 Robustness to the Number of Social Biases
We investigate the effectiveness of our proposed method under continual debiasing settings. We perform FAST on different knowledge sets in sequence and evaluate their performance. Results are reported in Table 23. It can be observed that SS obtained in the front stages is steady across the following stages, which indicates that after calibration on other knowledge, existing stored knowledge is still retained. Besides, LMS and ICAT even increase slightly in the process. These results prove the feasibility of continually updating the perception within language models.
@@ Appendix E Limitation and Future Works
While our research yields important contributions, we acknowledge the presence of certain limitations. Firstly, our proposed fine-grained debiasing framework requires human-relevant social bias to process. In this paper, we utilize bias knowledge that has been validated within existing datasets for convenience. In practice, retaining a comprehensive bias knowledge base is both time-consuming and labor-intensive. We notice that recent works Sahoo et al. (2022); Dev et al. (2023) have proposed an automated social bias detection method. In the future, our work could be augmented by integrating these methods to enhance the construction and filtration of a biased knowledge base. Besides, social bias in open language generation or dialogue Yu et al. (2022); Ovalle et al. (2023) represents another critical scenario for applying mitigating techniques, which is not addressed in this paper. Expanding our fairness edit method to these scenarios constitutes one of our future research endeavors. Finally, compared to the results on BERT and GPT2, the debiasing performance on larger models (Section 4.2) appears less pronounced. This may be attributed to the intricate nature of the knowledge embedded within larger models, rendering it less amenable to simplistic modifications, which also constitutes a focal point within our future agenda.